# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ A.9. Not Found / #404 (2023)
Figure 12. Not Found #404 on Opensea
@@
Not Found #404 is a 1-of-1 NFT forged “in memory of the absent.” that Pak created in August 2023 as a poignant tribute to the late crypto artist Alotta Money. The concept revolves around absence as presence. It leverages the web’s familiar “404 Not Found” error – the message one sees when a webpage is missing – as an artistic statement. Pak minted 404 with no metadata: no image, no name (beyond the token ID), no description . In Pak’s words, “#404 is a token forged in memory of the absent… Void of metadata by design to amplify the presence of absence.” The NFT exists as a sort of intentional glitch or blank, symbolizing the void left by Alotta Money’s passing. Socially, it speaks to how loss is felt in the digital age – even on a blockchain, an empty space can carry profound meaning as a memorial.
@@
Technically, 404 is engineered to be an NFT that always produces a 404 error for its metadata endpoint . Typically, an NFT’s metadata link provides JSON data including the name, description, and image URI. Pak deliberately set up 404 so that when any application (like OpenSea or Etherscan) queries the token’s metadata, the server responds with a “Not Found” status. This means the token has no visible content – no name, no image – just its token ID (#404 on that contract) . However, embedded within the smart contract is a tiny message: instead of the usual metadata, Pak hard-coded a plain string that says “In memory of the absent” as the response . That phrase is the only “content” of the NFT, acting like a eulogy delivered in code. The dynamic effect is that on any NFT marketplace, 404 appears as an empty or broken entry (which is itself conspicuous). Over time, assuming the piece stays as is, it will forever display as a missing asset, effectively eternalizing the idea of absence. This is a case where the lack of data is the feature, not a bug – a radical use of the NFT format to signify zero.
@@
404 was part of a charity auction (the Alotta Money Memorial on MakersPlace), so the direct participation was in the bidding process. Collectors vied not for a flashy image but for the honor of owning this conceptual piece. The winning participant, prominent collector WhaleShark, essentially paid a large sum for an NFT that “looks” like a broken link. In a broader sense, the audience participation comes in the form of interpretation: everyone who views the token on marketplaces just sees an error or blank, and must mentally fill in what it represents. Interestingly, because the contract ensures the metadata always returns a 404 error , any platform displaying it will show a broken image icon or “Not Found” text, inviting each viewer to participate by remembering or inquiring about Alotta Money and the context. Thus, while not interactive in a traditional way, it engages the audience’s awareness and curiosity. The community also participated in spreading the story – tweets and articles explained what 404 meant, effectively crowdsourcing the task of giving this “silent” NFT a voice.
@@
Pak had meticulous control over 404’s presentation. By deciding to include no metadata and making the token unchangeable (immutable after mint), Pak ensures that neither he nor the owner can ever add an image or title later. This permanence is part of the tribute’s integrity. The artist also controlled the context: 404 was introduced with a tweet explaining its purpose and dedication , guiding the initial audience understanding. After the auction, however, Pak’s control yielded to the token’s behavior in the wild – any confusion or discussion arising from people seeing a blank NFT became part of the piece’s life. It’s worth noting that creating an NFT that defies normal display could have negative repercussions (some might think it’s an error), but Pak accepted that risk to preserve the concept. In terms of curation, Pak placed 404 in a charity event so that its sale (48 ETH) benefited a cause (Alotta’s cancer fund or a related charity), aligning control of proceeds with the homage. Ultimately, Pak’s control was about ensuring emptiness – an ironic but powerful form of control where doing “nothing” was a deliberate artistic act.
Collective Emergent Behaviour
Once 404 was out in the world, the NFT community responded collectively by imbuing it with significance. Discussion threads emerged about “the NFT with no data” and what it meant; in this way, the community co-authored the narrative by sharing the backstory. Some artists were inspired by this approach and contemplated similar “negative space” works. As for the owner, WhaleShark displayed 404 in virtual galleries and on social media, where the emptiness became thought-provoking. Rather than diminishing interest, the lack of an image actually drew people in – a reverse of the usual NFT hype cycle. A subtle emergent behavior is that marketplaces had to handle a token with no metadata gracefully; some updated their interface to show “Unnamed” or simply the token ID. This sparked minor technical conversations in developer circles about how to index such an NFT, thereby 404 gently pushed the boundaries of NFT platform expectations. The most heartening emergent behavior was the collective act of remembrance: the NFT community, normally fixated on visuals and rarity, paused to remember an artist through a non-visual artifact. In that sense, 404 succeeded in creating a communal moment of silence, as it were, in the bustling NFT space.
@@
404 is pure poetry in digital form. Its poetry lies in what is absent: it forces the viewer to confront a void and find meaning in it. Much like John Cage’s silent composition “4’33”” or Rauschenberg’s erased drawing, Pak’s 404 finds art in nothingness, which in this context becomes a profound statement about loss. The use of the “Not Found” web error as the medium ties our sense of missing information to the emotion of missing a person. It’s a requiem encoded as a glitch. The token number 404 itself is meaningful – in web lore, 404 symbolizes a dead link, something that was once there but no longer accessible, mirroring Alotta Money’s departure. Moreover, 404 leverages the permanence of blockchain to eternalize an impermanent idea (absence). Its message “In memory of the absent” resonates universally – it’s not just about one individual, but about all those we’ve lost (it invites anyone to project their own feelings of loss onto it). As a social/artistic message, it reminds the tech-forward NFT community of human mortality and the gap that death leaves in our networks. In its quiet way, 404 perhaps also comments on the oversaturated NFT market – amid thousands of gaudy images, the most meaningful piece can be an empty one. The poetry of 404 lies in its silence and subtlety, making the presence of an absence palpably felt.
@@
404 was minted on 31 August 2023 and auctioned the next day at the “Alotta Money Tribute” event on MakersPlace . It sold for 48 ETH (roughly \$76,000 at the time) to collector WhaleShark , with 100% of the sale proceeds going to charity. The token resides on a custom contract; it is token ID 404 and notably the only token in that contract (no other token IDs exist, making 404 both the ID and the collection name) . Its on-chain metadata call returns a 404 error code by design , and the token carries a brief on-chain text “In memory of the absent.” The piece garnered considerable media attention for an NFT – features in art publications highlighted it as the first NFT that intentionally “does nothing” visually yet carries deep meaning. As of 2025, WhaleShark has not listed 404 for resale (and likely never will), aligning with its status as a memorial piece. 404 stands as Pak’s first major NFT after Merge, and while it’s a single edition, its impact was amplified by the hundreds of artists and collectors who witnessed and shared in the commemorative moment it represents.

### v2

@@ A.9. Not Found / #404 (2023)
Refer to caption A screenshot of the OpenSea collection page for “Not Found” by Pak. The header shows the collection name with a verified badge, labeled “BY PAK,” “ETHEREUM,” “1 item,” and “SEP 2023.” Below, the single NFT entry displays only a generic broken-image placeholder icon (a small mountain landscape silhouette inside a square) on a dark background, labeled “number 404, rank 1, Not listed.” No artwork image, title, or description is shown.
Figure 13. Not Found #404 on OpenSea (2023). The token was minted with no metadata by design—no image, no name, no description—so that any platform querying it receives only a 404 error. The resulting broken-image placeholder becomes the artwork itself: a deliberate void memorializing the absent.
@@
Not Found #404 is a 1-of-1 NFT forged “in memory of the absent” that Pak created in August 2023 as a poignant tribute to the late crypto artist Alotta Money. The concept revolves around absence as presence. It leverages the web’s familiar “404 Not Found” error—the message one sees when a webpage is missing—as an artistic statement. Pak minted 404 with no metadata: no image, no name (beyond the token ID), no description. In Pak’s words, “#404 is a token forged in memory of the absent… Void of metadata by design to amplify the presence of absence.” The NFT exists as a sort of intentional glitch or blank, symbolizing the void left by Alotta Money’s passing. Socially, it speaks to how loss is felt in the digital age—even on a blockchain, an empty space can carry profound meaning as a memorial.
@@
Technically, 404 is engineered to be an NFT that always produces a 404 error for its metadata endpoint. Typically, an NFT’s metadata link provides JSON data including the name, description, and image URI. Pak deliberately set up 404 so that when any application (like OpenSea or Etherscan) queries the token’s metadata, the server responds with a “Not Found” status. This means the token has no visible content—no name, no image—just its token ID (#404 on that contract). However, embedded within the smart contract is a tiny message: instead of the usual metadata, Pak hard-coded a plain string that says “In memory of the absent” as the response. That phrase is the only “content” of the NFT, acting like a eulogy delivered in code. The dynamic effect is that on any NFT marketplace, 404 appears as an empty or broken entry (which is itself conspicuous). Over time, assuming the piece stays as is, it will forever display as a missing asset, effectively eternalizing the idea of absence. This is a case where the lack of data is the feature, not a bug—a radical use of the NFT format to signify zero.
@@
404 was part of a charity auction (the Alotta Money Memorial on MakersPlace), so the direct participation was in the bidding process. Collectors vied not for a flashy image but for the honor of owning this conceptual piece. The winning participant, prominent collector WhaleShark, essentially paid a large sum for an NFT that “looks” like a broken link. In a broader sense, the audience participation comes in the form of interpretation: everyone who views the token on marketplaces just sees an error or blank, and must mentally fill in what it represents. Interestingly, because the contract ensures the metadata always returns a 404 error, any platform displaying it will show a broken image icon or “Not Found” text, inviting each viewer to participate by remembering or inquiring about Alotta Money and the context. Thus, while not interactive in a traditional way, it engages the audience’s awareness and curiosity. The community also participated in spreading the story—tweets and articles explained what 404 meant, effectively crowdsourcing the task of giving this “silent” NFT a voice.
@@
Pak had meticulous control over 404’s presentation. By deciding to include no metadata and making the token unchangeable (immutable after mint), Pak ensures that neither Pak nor the owner can ever add an image or title later. This permanence is part of the tribute’s integrity. The artist also controlled the context: 404 was introduced with a tweet explaining its purpose and dedication, guiding the initial audience understanding. After the auction, however, Pak’s control yielded to the token’s behavior in the wild—any confusion or discussion arising from people seeing a blank NFT became part of the piece’s life. It’s worth noting that creating an NFT that defies normal display could have negative repercussions (some might think it’s an error), but Pak accepted that risk to preserve the concept. In terms of curation, Pak placed 404 in a charity event so that its sale (48 ETH) benefited a cause (Alotta’s cancer fund or a related charity), aligning control of proceeds with the homage. Ultimately, Pak’s control was about ensuring emptiness—an ironic but powerful form of control where doing “nothing” was a deliberate artistic act.
Collective Emergent Behavior
Once 404 was out in the world, the NFT community responded collectively by imbuing it with significance. Discussion threads emerged about “the NFT with no data” and what it meant; in this way, the community co-authored the narrative by sharing the backstory. Some artists were inspired by this approach and contemplated similar “negative space” works. As for the owner, WhaleShark displayed 404 in virtual galleries and on social media, where the emptiness became thought-provoking. Rather than diminishing interest, the lack of an image actually drew people in—a reverse of the usual NFT hype cycle. A subtle emergent behavior is that marketplaces had to handle a token with no metadata gracefully; some updated their interface to show “Unnamed” or simply the token ID. This sparked minor technical conversations in developer circles about how to index such an NFT, thereby 404 gently pushed the boundaries of NFT platform expectations. The most heartening emergent behavior was the collective act of remembrance: the NFT community, normally fixated on visuals and rarity, paused to remember an artist through a non-visual artifact. In that sense, 404 succeeded in creating a communal moment of silence, as it were, in the bustling NFT space.
@@
404 is pure poetry in digital form. Its poetry lies in what is absent: it forces the viewer to confront a void and find meaning in it. Much like John Cage’s silent composition “4’33”” or Rauschenberg’s erased drawing, Pak’s 404 finds art in nothingness, which in this context becomes a profound statement about loss. The use of the “Not Found” web error as the medium ties our sense of missing information to the emotion of missing a person. It’s a requiem encoded as a glitch. The token number 404 itself is meaningful—in web lore, 404 symbolizes a dead link, something that was once there but no longer accessible, mirroring Alotta Money’s departure. Moreover, 404 leverages the permanence of blockchain to eternalize an impermanent idea (absence). Its message “In memory of the absent” resonates universally—it’s not just about one individual, but about all those we’ve lost (it invites anyone to project their own feelings of loss onto it). As a social/artistic message, it reminds the tech-forward NFT community of human mortality and the gap that death leaves in our networks. In its quiet way, 404 perhaps also comments on the oversaturated NFT market—amid thousands of gaudy images, the most meaningful piece can be an empty one. The poetry of 404 lies in its silence and subtlety, making the presence of an absence palpably felt.
@@
404 was minted on 31 August 2023 and auctioned the next day at the “Alotta Money Tribute” event on MakersPlace. It sold for 48 ETH (roughly \$78,000 at the time) to collector WhaleShark, with sale proceeds going to charity. The token resides on a custom contract; it is token ID 404 and notably the only token in that contract (no other token IDs exist, making 404 both the ID and the collection name). Its on-chain metadata call returns a 404 error code by design, and the token carries a brief on-chain text “In memory of the absent.” The piece garnered considerable media attention for an NFT—features in art publications highlighted it as the first NFT that intentionally “does nothing” visually yet carries deep meaning. As of 2025, WhaleShark has not listed 404 for resale (and likely never will), aligning with its status as a memorial piece. 404 stands as one of Pak’s later protocol artworks, and while it’s a single edition, its impact was amplified by the hundreds of artists and collectors who witnessed and shared in the commemorative moment it represents.
Annotation
Not Found is the portfolio’s limiting case: a work that strips protocol art to its conceptual minimum. C1 (system-centric composition) is expressed through pure negation—the artwork’s meaning resides entirely in the contract’s deliberate refusal to return metadata, making the absence of content the content itself. C6 (poetic message embedding) is correspondingly intense: the token ID 404 and the lone on-chain string “In memory of the absent” compress an entire elegy into the structure of a failed HTTP request. Unlike every other work in the portfolio, Not Found involves no C3 (distributed agency), no C5 (economy-driven engagement beyond the charity auction), and no C7 (composability). Its isolation is the point: it demonstrates that protocol art need not rely on mass participation or elaborate tokenomics to achieve poetic force—a single, precisely engineered absence can suffice.